# Sufficient and Necessary Barrier-like Conditions for Safety and Reach-avoid Verification of Stochastic Discrete-time Systems

### v2

@@ IV-B Reach-avoid Verification II
Let $\epsilon_{2}\in[0,1)$. Under Assumption 2, there exist a constant $\gamma\in(0,1)$ and a function $v(\bm{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$, which is bounded over $\mathcal{X}$ and satisfies the following condition:
@@
if and only if $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})\geq\epsilon_{2}$.
@@
1) We first prove the “only if” part.
Since $v(\bm{x})$ satisfies (14), we can obtain, following the induction of (13) by replacing “=” with $\leq$”, that
$\begin{split}v(\bm{x})\leq\tilde{V}_{\gamma}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{i\rightarrow\infty}h_{i}(\bm{x})\end{split}$
for $\bm{x}\in\mathbb{R}^{n}$, where $\tilde{V}_{\gamma}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is the value function defined in (11). Since $v(\bm{\phi}_{\pi}^{\bm{x}}(i+1))\leq 0$ when $\bm{\phi}_{\pi}^{\bm{x}}(i+1)\in\mathbb{R}^{n}\setminus\mathcal{X}$, we have
$\begin{split}h_{i}(\bm{x})&=\gamma^{i+1}\mathbb{E}_{\pi}[\prod_{j=1}^{i}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))v(\bm{\phi}_{\pi}^{\bm{x}}(i+1))]\leq\gamma^{i+1}\mathbb{E}_{\pi}[\prod_{j=1}^{i}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))w_{i+1}(\bm{x})],\end{split}$
where $w_{i+1}(\bm{x})=1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(i+1))v(\bm{\phi}_{\pi}^{\bm{x}}(i+1))$. Also, since $v(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is bounded over $\mathcal{X}$ and $\lim_{i\rightarrow\infty}\gamma^{i+1}=0$, we conclude
$\lim_{i\rightarrow\infty}h_{i}(\bm{x})=0$
@@
2) We will prove the “if” part.
According to Lemma 3, $\lim_{\gamma\rightarrow 1^{-}}\tilde{V}_{\gamma}(\bm{x}_{0})=\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$ holds. Since $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})>\epsilon_{2}$, there exists $\gamma_{0}$ such that $\tilde{V}_{\gamma_{0}}(\bm{x}_{0})\geq\epsilon_{2}$ according to Lemma 3. Moreover, according to Proposition 5, $\tilde{V}_{\gamma_{0}}(\bm{x})$ satisfies
@@
If we consider an initial set $\mathcal{X}_{0}\subseteq\mathcal{X}\setminus\mathcal{X}_{r}$, which includes infinite initial states, rather than a fixed initial state $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$, we cannot guarantee that there exist a constant $\gamma\in(0,1)$ and a function $v(\bm{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$, which is bounded over $\mathcal{X}$ and satisfies the condition (14) with $v(\bm{x})\geq\epsilon_{2},\forall\bm{x}\in\mathcal{X}_{0}$ replacing $v(\bm{x}_{0})\geq\epsilon_{2}$, such that $\mathbb{P}_{\pi}(RA_{\bm{x}})\geq\epsilon_{2},\forall\bm{x}\in\mathcal{X}_{0}$. This is because we cannot guarantee that $\lim_{\gamma\rightarrow 1^{-}}\tilde{V}_{\gamma}(\bm{x})=\mathbb{P}_{\pi}(RA_{\bm{x}})$ holds uniformly over $\mathcal{X}_{0}$.
In addition, condition (14) is a typical instance of condition (13) with $\alpha>1$ and $\beta=0$ in Theorem 5 in [26], which offers lower bounds of the reach-avoid probability in the context of finite-time reach-avoid verification. $\blacksquare$
@@
It is worth noting here that we can also construct a necessary and sufficient condition to certify upper bounds of the safety probability $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X})$ such that the system (1) starting from the initial state $\bm{x}_{0}$ will stay within the safe set $\mathcal{X}$ for all time [29], under the assumption that $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X})<1-\epsilon_{1}$. Under the assumption that $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X})<1-\epsilon_{1}$, there exist a constant $\gamma\in(0,1)$ and a function $v(\bm{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$, which is bounded over $\mathcal{X}$ and satisfies the following condition:
@@
if and only if $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X})\leq 1-\epsilon_{1}$ (or equivalently, $\mathbb{P}_{\pi}(\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}\in\mathbb{R}^{n}\setminus\mathcal{X})\geq\epsilon_{1}$). Condition (15) is also a typical instance of condition (6) with $\alpha>1$ and $\beta=0$ in Theorem 2 in [26], which offers upper bounds of the safety probability in the finite-time safety verification. Such conditions for certifying upper bounds become particularly significant in scenarios where $\mathcal{R}^{n}\setminus\mathcal{X}$ represents the target set that the system aims to reach. In this context, the safety probability will be referred to as the liveness probability. $\blacksquare$
Based on the value function (11), we are able to show the necessity of another sufficient barrier-like condition in [27] for the reach-avoid verification under Assumption 2. The condition is presented below:
@@
where $\Omega$ is a set in (5). If there exist a function $v(\cdot):\Omega\rightarrow\mathbb{R}$ and a bounded function $w(\cdot):\Omega\rightarrow\mathbb{R}$ satisfying (16), $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})\geq\epsilon_{2}$ holds. This conclusion can be justified by following the proof of Corollary 2 in [27]. In the following, we just demonstrate its necessity.
@@
$\begin{split}\frac{\mathbb{E}_{\bm{\theta}}[w(\bm{f}(\bm{x},\bm{\theta}))]-w(\bm{x})-\tilde{V}_{\gamma_{0}}(\bm{x})}{1+\gamma_{1}}=&\frac{\gamma_{1}\mathbb{E}_{\bm{\theta}}[\tilde{V}_{\gamma_{0}}(\bm{f}(\bm{x},\bm{\theta}))]-\gamma_{1}\tilde{V}_{\gamma_{0}}(\bm{x})-\tilde{V}_{\gamma_{0}}(\bm{x})}{1+\gamma_{1}}\\ =&\frac{\gamma_{1}}{1+\gamma_{1}}\mathbb{E}_{\bm{\theta}}[\tilde{V}_{\gamma_{0}}(\bm{f}(\bm{x},\bm{\theta}))]-\tilde{V}_{\gamma_{0}}(\bm{x})\\ \geq&\gamma_{0}\mathbb{E}_{\bm{\theta}}[\tilde{V}_{\gamma_{0}}(\bm{f}(\bm{x},\bm{\theta}))]-\tilde{V}_{\gamma_{0}}(\bm{x})=0.\end{split}$
@@
V Conclusion

### v3

@@ IV-B Reach-avoid Verification II
Let $\epsilon_{2}\in[0,1)$. If there exist a constant $\gamma\in(0,1)$ and a function $v(\bm{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$, which is bounded over $\mathcal{X}$ and satisfies the following condition:
@@
then, $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})\geq\epsilon_{2}$. Moreover, under Assumption 2, there indeed exist such a constant $\gamma\in(0,1)$ and a function $v(\bm{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$, bounded over $\mathcal{X}$, that satisfy (13).
@@
1) Since $v(\bm{x})$ satisfies (13), by following the inductive argument used in the proof of Proposition 4-where we showed
$v(\bm{x})=\tilde{V}_{\gamma}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{k\to\infty}h_{k}(\bm{x})$
-but replacing the equality “=” with $\leq$”, we obtain that
$\begin{split}v(\bm{x})\leq\tilde{V}_{\gamma}(\bm{x})+1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{x})\lim_{k\rightarrow\infty}h_{k}(\bm{x})\end{split}$
for $\bm{x}\in\mathbb{R}^{n}$, where $\tilde{V}_{\gamma}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is the value function defined in (11). Since $v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))\leq 0$ when $\bm{\phi}_{\pi}^{\bm{x}}(k+1)\in\mathbb{R}^{n}\setminus\mathcal{X}$, we have
$\begin{split}h_{k}(\bm{x})&=\gamma^{k+1}\mathbb{E}_{\pi}[\prod_{j=1}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))]\leq\gamma^{k+1}\mathbb{E}_{\pi}[\prod_{j=1}^{k}1_{\mathcal{X}\setminus\mathcal{X}_{r}}(\bm{\phi}_{\pi}^{\bm{x}}(j))w_{k+1}(\bm{x})],\end{split}$
where $w_{k+1}(\bm{x})=1_{\mathcal{X}}(\bm{\phi}_{\pi}^{\bm{x}}(k+1))v(\bm{\phi}_{\pi}^{\bm{x}}(k+1))$. Also, since $v(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is bounded over $\mathcal{X}$ and $\lim_{k\rightarrow\infty}\gamma^{k+1}=0$, we conclude
$\lim_{k\rightarrow\infty}h_{k}(\bm{x})=0$
@@
2) According to Lemma 3, $\lim_{\gamma\rightarrow 1^{-}}\tilde{V}_{\gamma}(\bm{x}_{0})=\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})$ holds. Since $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})>\epsilon_{2}$, there exists $\gamma_{0}$ such that $\tilde{V}_{\gamma_{0}}(\bm{x}_{0})\geq\epsilon_{2}$ according to Lemma 3. Moreover, according to Proposition 5, $\tilde{V}_{\gamma_{0}}(\bm{x})$ satisfies
@@
If we consider an initial set $\mathcal{X}_{0}\subseteq\mathcal{X}\setminus\mathcal{X}_{r}$, which includes infinitely many initial states, rather than a fixed initial state $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$, we cannot guarantee that there exist a constant $\gamma\in(0,1)$ and a function $v(\bm{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$, which is bounded over $\mathcal{X}$ and satisfies the condition (13) with $v(\bm{x})\geq\epsilon_{2},\forall\bm{x}\in\mathcal{X}_{0}$ replacing $v(\bm{x}_{0})\geq\epsilon_{2}$, such that $\mathbb{P}_{\pi}(RA_{\bm{x}})\geq\epsilon_{2},\forall\bm{x}\in\mathcal{X}_{0}$. This is because we cannot guarantee that $\lim_{\gamma\rightarrow 1^{-}}\tilde{V}_{\gamma}(\bm{x})=\mathbb{P}_{\pi}(RA_{\bm{x}})$ holds uniformly over $\mathcal{X}_{0}$.
In addition, condition (13) is a typical instance of condition (13) with $\alpha>1$ and $\beta=0$ in Theorem 5 in [xue2024finite], which offers lower bounds of the reach-avoid probability in the context of finite-time reach-avoid verification. $\blacksquare$
@@
We note here that we can also construct a necessary and sufficient condition to certify upper bounds of the safety probability $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X})$ such that the system (1) starting from the initial state $\bm{x}_{0}$ will stay within the safe set $\mathcal{X}$ for all time [yu2023safe], under the assumption that $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X})<1-\epsilon_{1}$. Under the assumption that $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X})<1-\epsilon_{1}$, there exist a constant $\gamma\in(0,1)$ and a function $v(\bm{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$, which is bounded over $\mathcal{X}$ and satisfies the following condition:
@@
if and only if $\mathbb{P}_{\pi}(\forall k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}(k)\in\mathcal{X})\leq 1-\epsilon_{1}$ (or equivalently, $\mathbb{P}_{\pi}(\exists k\in\mathbb{N}.\bm{\phi}^{\bm{x}_{0}}_{\pi}\in\mathbb{R}^{n}\setminus\mathcal{X})\geq\epsilon_{1}$). Condition (14) is also a typical instance of condition (6) with $\alpha>1$ and $\beta=0$ in Theorem 2 in [xue2024finite], which offers upper bounds of the safety probability in the finite-time safety verification. Such conditions for certifying upper bounds become particularly significant in scenarios where $\mathbb{R}^{n}\setminus\mathcal{X}$ represents the target set that the system aims to reach. In this context, the safety probability will be referred to as the liveness probability. $\blacksquare$
Based on the value function (11), we are able to show the necessity of another sufficient barrier-like condition in [xue2021reach] for the reach-avoid verification under Assumption 2. The condition is presented below:
@@
where $\Omega$ is a set in (5). If there exist a function $v(\cdot):\Omega\rightarrow\mathbb{R}$ and a bounded function $w(\cdot):\Omega\rightarrow\mathbb{R}$ satisfying (15), $\mathbb{P}_{\pi}(RA_{\bm{x}_{0}})\geq\epsilon_{2}$ holds. This conclusion can be justified by following the proof of Corollary 2 in [xue2021reach]. In the following, we just demonstrate its necessity.
@@
$\begin{split}&\frac{\mathbb{E}_{\bm{\theta}}[w(\bm{f}(\bm{x},\bm{\theta}))]-w(\bm{x})-\tilde{V}_{\gamma_{0}}(\bm{x})}{1+\gamma_{1}}\\ =&\frac{\gamma_{1}\mathbb{E}_{\bm{\theta}}[\tilde{V}_{\gamma_{0}}(\bm{f}(\bm{x},\bm{\theta}))]-\gamma_{1}\tilde{V}_{\gamma_{0}}(\bm{x})-\tilde{V}_{\gamma_{0}}(\bm{x})}{1+\gamma_{1}}\\ =&\frac{\gamma_{1}}{1+\gamma_{1}}\mathbb{E}_{\bm{\theta}}[\tilde{V}_{\gamma_{0}}(\bm{f}(\bm{x},\bm{\theta}))]-\tilde{V}_{\gamma_{0}}(\bm{x})\\ \geq&\gamma_{0}\mathbb{E}_{\bm{\theta}}[\tilde{V}_{\gamma_{0}}(\bm{f}(\bm{x},\bm{\theta}))]-\tilde{V}_{\gamma_{0}}(\bm{x})=0.\end{split}$
@@
V Examples
In this section, we demonstrate the application of our theoretical developments through two examples. In both cases, the function $f(x,\theta)$ is a polynomial in the state variables $x$, and the safe set $\mathcal{X}$ as well as the target set $\mathcal{X}_{r}$ are semi-algebraic sets. We aim to search for polynomial barrier-like functions to solve the associated verification problem. To do this, we encode the constraints (4), (10), and (13) as semi-definite programs (SDPs) using the sum of squares (SOS) decomposition for multivariate polynomials. The resulting SDPs are then solved the tool Mosek 10.1.21 [aps2019mosek]. To ensure numerical stability during the solution of these SDPs, we impose a constraint on the coefficients of the unknown polynomials, specifically restricting them to the interval [100,100][-100,100]. In the sequel, $\sum[x]$ denotes the set of sum-of-squares polynomials over variables $x$, i.e., $\sum[x]=\{p\in\mathbb{R}[x]\mid p=\sum_{i=1}^{k}q^{2}_{i}(x),q_{i}(x)\in\mathbb{R}[x],i=1,\ldots,k\}$, where $\mathbb{R}[x]$ denotes the ring of polynomials in variables $x$.
Example 1 (Safety Verification)
Consider the one-dimensional discrete-time system:
$x(l+1)=(-0.5+\theta(l))\,x(l),$ (16)
where $\theta(l)\in\Theta=[-1,1]$ is uniform, the safe set is $\mathcal{X}=\{x\mid h(x)\leq 0\}$ with $h(x)=x^{2}-1$, and the initial state is $x_{0}=-0.8$. We simulate $10^{4}$ trajectories over $10^{4}$ time steps. The estimated safety probability is $0.8286$. Figure 1 shows three example trajectories over 10 time steps.
Figure 1: Trajectories of system in Example 1 with $x_{0}=-0.8$.The region enclosed by the red curve represents the safe set $\mathcal{X}$, and the blue points correspond to the system states visited over 10 time steps.
SDP formulation: We solve the safety verification problem with $\epsilon_{1}=0.65$ and $\epsilon_{1}=0.75$, as defined in Definition 3, via solving the constraint (4). The corresponding SDP over unknown polynomials $(v(x),s_{0}(x),s_{1}(x))$ is:
$\begin{cases}1-\epsilon_{1}-v(x_{0})\geq 0,\\ v(x)-\mathbb{E}_{\theta}[v(f(x,\theta))]+s_{0}(x)h(x)\in\sum[x],\\ v(x)-1-s_{1}(x)h(x)\in\sum[x],\\ v(x)\in\sum[x],\quad s_{0}(x),s_{1}(x)\in\sum[x].\end{cases}$
[…]
TABLE I: SDP feasibility for Example 1 with $x_{0}=-0.8$
(✔: feasible; ✗: infeasible)
Degree $\epsilon_{1}=0.65$ $\epsilon_{1}=0.75$
6
18
Extension to initial set (as indicated in Remark 1): For initial set $\mathcal{X}_{0}=\{x\mid h_{0}(x)\leq 0\}$ with $h_{0}(x)=(x-0.8)^{2}-0.01$, the resulting SDP over unknown polynomials $(v(x),s_{i}(x),i=0,\ldots,2)$ is
$\begin{cases}1-\epsilon_{1}-v(x)+s_{0}(x)h_{0}(x)\in\sum[x],\\ v(x)-\mathbb{E}_{\theta}[v(f(x,\theta))]+s_{1}(x)h(x)\in\sum[x],\\ v(x)-1-s_{2}(x)h(x)\in\sum[x],\\ v(x)\in\sum[x],s_{0}(x)\in\sum[x],\\ s_{1}(x)\in\sum[x],s_{2}(x)\in\sum[x].\end{cases}$
It is feasible for degree 18 but infeasible for degree 16, as reported in Table II. We also compute an empirical estimate of the safety probability using a Monte Carlo method. Specifically, we draw $10^{3}$ initial states independently from the initial set $\mathcal{X}_{0}$ according to a uniform distribution, and for each initial state, we simulate $10^{4}$ trajectories over $10^{4}$ time steps. This procedure yields an estimated safety probability of $0.7521$.
TABLE II: SDP feasibility for Example 1 with the initial set $\mathcal{X}_{0}$
(✔: feasible; ✗: infeasible)
Degree $\epsilon_{1}=0.60$ $\epsilon_{1}=0.65$
14
22
Example 2 (Reach-Avoid Verification)
Consider the same system as Example 1. The safe set is $\mathcal{X}=\{x\mid h(x)\leq 0\}$ with $h(x)=x^{2}-1$, the target set is $\mathcal{X}_{r}=\{x\mid g(x)\leq 0\}$ with $g(x)=x^{2}-0.01$, and the initial state is $x_{0}=-0.8$. We simulate $10^{4}$ trajectories over $10^{4}$ time steps. The estimated reach-avoid probability is $0.8240$. Figure 2 shows three example trajectories over 10 time steps.
Figure 2: Trajectories of the system in Example 2 starting from $x_{0}=-0.8$. The regions enclosed by the red and green curves represent the safe set $\mathcal{X}$ and the target set $\mathcal{X}_{r}$, respectively, while the blue points indicate the system states visited over 10 time steps.
SDP formulation: We solve the reach-avoid verification problem with $\epsilon_{2}=0.65$ and $\epsilon_{2}=0.75$, as defined in Definition 4, by solving the constraints (10) and (13). As proven in Proposition 6 (see Appendix), for any $x\in\mathcal{X}\setminus\mathcal{X}_{r}$, the trajectory will leave this set in finite time with probability one. Therefore, Assumption 1 is satisfied. Therefore, we can address the reach-avoid verification problem by solving constraint (10). The key difference between constraints (10) and (13) lies in the treatment of the term $\mathbb{E}_{\theta}[v(f(x,\theta))]$: in (10), this term is multiplied by $1$, whereas in (13), it is scaled by a discount factor $\gamma\in(0,1)$. As a result, their corresponding SDPs—formulated over the unknown polynomials ($v(x)$, $s_{i}(x)$, $i=0,\ldots,3$)—can be expressed in a unified form by treating $\gamma$ as a tunable parameter, as shown below. Specifically, setting $\gamma=1$ recovers constraint (10), while choosing $\gamma\in(0,1)$ corresponds to constraint (13).
$\begin{cases}v(x_{0})-\epsilon_{2}\geq 0,\\ \gamma\mathbb{E}_{\theta}[v(f(x,\theta))]-v(x)+s_{0}(x)h(x)-s_{1}(x)g(x)\in\sum[x],\\ 1-v(x)+s_{2}(x)g(x)\in\sum[x],\\ -v(x)-s_{3}(x)h(x)\in\sum[x],\\ s_{0}(x)\in\sum[x],s_{1}(x)\in\sum[x],\\ s_{2}(x)\in\sum[x],s_{3}(x)\in\sum[x].\end{cases}$
SDP feasibility vs polynomial degree ($v(x)$, $s_{i}(x)$, $i=0,\ldots,3$) and $\gamma$: Table III summarizes feasibility results.
TABLE III: SDP feasibility for Example 2 with $\bm{x}_{0}=-0.8$
(✔: feasible; ✗: infeasible)
Degree $\epsilon_{2}$ $\gamma=1$ $\gamma=0.999$ $\gamma=0.99$
6 0.65
6 0.75
18 0.65
18 0.75
Extension to initial set (as indicated in Remark 4): For $\mathcal{X}_{0}=\{x\mid h_{0}(x)\leq 0\}$ with $h_{0}(x)=(x-0.8)^{2}-0.01$, the resulting SDP over unknown polynomials $(v(x),s_{i}(x),i=0,\ldots,4)$ is
$\begin{cases}v(x)-\epsilon_{2}+s_{0}(x)h_{0}(x)\in\sum[x],\\ \gamma\mathbb{E}_{\theta}[v(f(x,\theta))]-v(x)+s_{1}(x)h(x)-s_{2}(x)g(x)\in\sum[x],\\ 1-v(x)+s_{3}(x)g(x)\in\sum[x],\\ -v(x)-s_{4}(x)h(x)\in\sum[x],\\ s_{0}(x)\in\sum[x],s_{1}(x)\in\sum[x],\\ s_{2}(x)\in\sum[x],s_{3}(x)\in\sum[x],s_{4}(x)\in\sum[x].\end{cases}$
Its feasibility summarized in Table IV. We also compute an empirical estimate of the reach-avoid probability using a Monte Carlo method. Specifically, we draw $10^{3}$ initial states independently from the initial set $\mathcal{X}_{0}$ according to a uniform distribution, and for each initial state, we simulate $10^{4}$ trajectories over $10^{4}$ time steps. This procedure yields an estimated reach-avoid probability of $0.7510$.
TABLE IV: SDP feasibility for Example 2 with the initial set $\mathcal{X}_{0}$
(✔: feasible; ✗: infeasible)
Degree $\epsilon_{2}$ $\gamma=1$ $\gamma=0.999$ $\gamma=0.99$
20 0.65
22 0.65
The above examples demonstrate how the choice of polynomial degree and the factor $\gamma$ affect the feasibility of the safety and reach–avoid verification problems. Increasing the polynomial degree enhances the representational capacity of the barrier-like functions, which benefits both safety and reach–avoid verification by enabling the SDP to satisfy the associated conditions for larger tolerance parameters $\epsilon_{1}$ and $\epsilon_{2}$. In contrast, the discount factor $\gamma$, which is specific to the reach-avoid verification, influences the trade-off between conservatism and feasibility: values of $\gamma$ closer to 1 generally make the SDP more likely to be feasible under less conservative conditions.
VI Conclusion